Cassie has 33 coins in her purse consisting of 5-peso coins and 10- peso coins. The
number of 10 - peso coins is twice the number of 5 - peso coins. If the total amount in
the purse is PhP 275.00, How many 5 - peso coins are there? How many 10 - peso coins are there?

Answer by ikleyn(53765)

.

Let x be the number of the 5-peso coins.

Then the number of the 10-peso coins is  2x, according to the condition.


The total money equation is


    5x + 10*(2x) = 275.


Simplify and solve


    5x + 20x = 275

      25x    = 275

        x    = 275/25 = 11.


ANSWER.  11 of the 5-peso coins and 2*11 = 22  of the 10-peso cpins.


CHECK.   5*11 + 10*22 = 55 + 220 = 275 pesos.   ! Correct !

Solved.